OpenStudy (anonymous):

Write the equation y - 2 = 4(x + 5) in standard form.

OpenStudy (uri):

y-2=4x+20 y=4x+2+20 y=4x+22

OpenStudy (anonymous):

Standard form is ax + by= C. So, would that make the answer 4x + y=-22?

OpenStudy (uri):

Yup.

OpenStudy (anonymous):

Thanks so much! I needed to understand the order!!